How many grams of NO are required to produce 198 g of N2 in the following reaction?

4NH3(g) + 6NO (g) - 5N2(g) + 6H2O (l)
1
Expert's answer
2014-11-18T12:26:16-0500
We can use the method of proportion. The molar mass of 6NO = 180g/mole, the molar mass of 5N2 is 140g/mole. Suppose x is mass of NO. So, the proportion will be x/180 = 198g/140 andx = (180*198)/140 = 254.57g.